2NaOH+H2SO4 - 2H2O+Na2SO4
how many grams of sodium sulfate will be formed if you start with 250 grams of sodium hydroxide and you have an excess of sulfuric acid?
1
Expert's answer
2018-08-31T07:08:15-0400
First we must calculate the mole for NaOH.
n(NaOH)=250/40= 6.25 moles of NaOH.
From reaction we see than 2 moles of NaOH give 1 mole of Na2SO4. That means, that 6.25 moles of NaOH will give 6.25/2=3.125 moles of Na2SO4.
Then ee can calculate the mass of Na2SO4.
m(Na2SO4)= 3.125*142=443.75 g Na2SO4.